3. Suppose there is a Mendelian trait with incomplete dominance. Therefore, you can tell the genotype of every individual by its phenotype. If one of the two homozygotes occurs with a frequency of 6%, what would be the frequency in the population of the other homozygote?
Answer #1

Suppose there are two alleles, A and B.

AA = phenotype a

AB = phenotype b

BB = phenotype c

Genotypic frequency of AA = 6% = 6/100 = 0.06

Allelic frequency of A = √0.0 6 = 0.245

Therefore, allelic frequency of B = 1 - 0.245 = 0.755

Genotypic frequency of BB = 0.755 × 0.755 = 0.57 = 57%

Other homozygote would be 57%.

Please rate.